A led bullet of mass 0.05kg is fired with a velocity of 200m/s into a lead block of mass 0.95kg. Given that the lead block can move freely calculate the kinetic energy after the impact and the loss of energy


1
Expert's answer
2019-10-31T11:21:38-0400

From the conservation of energy:


"mV=(m+M)v"

"(0.05)(200)=(0.05+0.95)v"

"v=10\frac{m}{s}"

The kinetic energy after the impact:


"K=0.5(m+M)v^2=0.5(1)10^2=50\ J"

The loss of energy:


"E=0.5mV^2-K=0.5(0.05)(200)^2-50=950\ J"